OpenStudy (anonymous):

Write an equation for the line that contains the point (8, -3) and is parallel to the line x + 4y = 9. A. -4x + y = -35 B. x + 4y = -44 C. x + 4y = -20 D. x + 4y = -4

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tyteen4a03):

First trim down answers by plugging in numbers to see if LHS and RHS match, then calculate the slope of x + 4y = 9 and see if they match those in the answers.

OpenStudy (anonymous):

x+4y=9_(8,-3) 4y=-x+9 y=-(x)/(4)+(9)/(4) y=mx+b m=-(1)/(4) y=mx+b (-3)=(4)*(8)+b b=-35 y=4x-35